OpenStudy (anonymous):

Solve using the quadratic formula: x^2-3x-18=0

OpenStudy (anonymous):

It factors to: (x-6)(x+3)=0. So x=6,-3

OpenStudy (anonymous):

what about with the quadratic formula?

OpenStudy (anonymous):

Never mind!

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

I got the answer... I was adding incorrectly...

OpenStudy (anonymous):

ok